# GlotEval: A Test Suite for Massively Multilingual Evaluation of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are advancing at an unprecedented pace globally, with regions increasingly adopting these models for applications in their primary language. Evaluation of these models in diverse linguistic environments, especially in low-resource languages, has become a major challenge for academia and industry. Existing evaluation frameworks are disproportionately focused on English and a handful of high-resource languages, thereby overlooking the realistic performance of LLMs in multilingual and lower-resource scenarios. To address this gap, we introduce GlotEval, a lightweight framework designed for massively multilingual evaluation. Supporting seven key tasks (machine translation, text classification, summarization, open-ended generation, reading comprehension, sequence labeling, and intrinsic evaluation), spanning over dozens to hundreds of languages, GlotEval highlights consistent multilingual benchmarking, language-specific prompt templates, and non-English-centric machine translation. This enables a precise diagnosis of model strengths and weaknesses in diverse linguistic contexts. A multilingual translation case study demonstrates GlotEval’s applicability for multilingual and language-specific evaluations.
[…]
toolbox GlotEval: github.com/MaLA-LM/GlotEval
@@ 1 Introduction
In recent years, driven by rapid progress in natural language processing and deep learning, large language models (LLMs) such as GPT-4 (OpenAI, 2023) and Deepseek-R1 (DeepSeek-AI et al., 2025) have shown remarkable reasoning and generation capabilities across multiple languages and tasks. Although these models approach or surpass expert-level performance in certain high-resource languages (e.g., English), they often exhibit substantial performance fluctuations in other linguistic environments (Zhang et al., 2024). This discrepancy partially arises from the imbalance and scarcity of training data of low-source languages, and partially from the limited multilingual coverage of current evaluation frameworks: many were originally designed for English or a few widely spoken languages, making it difficult to extend them efficiently to more diverse linguistic tasks or to adapt custom prompts and configurations for each language. Meanwhile, as LLMs proliferate worldwide and different regions rely on their respective local languages, large-scale (massively) multilingual evaluation involving numerous low-resource languages has emerged as a critical research direction.
@@
    Consistent Multilingual Benchmarking. We integrated 20+ existing multilingual benchmarks into a unified pipeline, by standardizing all ISO 639-3 language codes in the different benchmarks  [^1] https://iso639-3.sil.org/about, which covers nearly all known languages worldwide. By aligning benchmark language identifiers with ISO 639-3 codes, we enable evaluations for specific languages or language groups (e.g., Bantu, Dravidian, or Uralic languages), allowing the framework to automatically search among integrated benchmarks to find matching test sets. This mapping also makes it easier to incorporate new large-scale benchmarks that target mid- or low-resource languages, ensuring flexibility for future expansions.
@@
    Language-Specific Prompt Templates. Users can configure prompts for each language individually, thereby enabling more precise assessments of a model’s instruction-following ability across diverse linguistic settings. All templates are maintained in a centralized prompt library that supports multilingual benchmarks, allowing easy customization as needed. In this way, each task within a benchmark can be run potentially using prompts in the task’s original language, rather than defaulting to English prompts. To simplify cross-lingual adaptation, we also implemented Microsoft Translator  [^2] https://learn.microsoft.com/en-us/azure/ai-services/translator/language-support integration that automatically propagates user-defined prompt templates from one single language to 130+ supported languages.
@@
    Non-English-Centered Machine Translation Evaluation. GlotEval is designed to break away from the traditional English-centric paradigm. Thanks to translation benchmarks featuring fully or partially multi-aligned datasets, GlotEval enables non-English-centered translation evaluations by allowing any supported language to serve as the pivot: users simply update the pivot language in the configuration, to assess “any-to-pivot” / “pivot-to-any” translation directions. This flexibility ensures that GlotEval breaks from the traditional “English $\leftrightarrow$ other language” paradigm and adapts seamlessly to diverse, potentially low-resource, language pairs.
By bringing all these capabilities together in a cohesive framework, GlotEval aims to facilitate large-scale, in-depth evaluations of multilingual LLMs across both widely spoken and underrepresented languages, ultimately driving forward more inclusive LLM evaluation. GlotEval can be a robust tool for researchers and developers conducting massively multilingual LLM evaluation.
@@ 2 Related Work
Several evaluation toolkits and benchmarks have been developed to systematically assess LLMs. EleutherAI’s LM Evaluation Harness (Gao et al., 2023) is a widely adopted framework covering over 60 tasks, including multilingual datasets such as XNLI (15 languages) and Belebele (122 languages). UltraEval (He et al., 2024) improves modularity and supports FLORES-200 for multilingual translation. OpenAI Evals [^3] https://github.com/openai/evals provides a highly flexible, community-driven framework, and OpenCompass (Contributors, 2023) offers a comprehensive platform with broad support for datasets and models. MEGA (Ahuja et al., 2023) evaluates generative LLMs across diverse languages, with a focus on standard NLP benchmarks. LightEval (Fourrier et al., 2023) developed a flexible LLM evaluation framework that supports different backends.
@@ 3.1 Benchmarks, Languages and Metrics
Tatoeba Tiedemann (2020) 500+ Mixed short sents. Yes (GitHub) BLEU, ChrF
OpenSubtitles v2024 Lison and Tiedemann (2016) 93 Subtitles Yes (Website) BLEU, ChrF
@@
Summarization XLSum Hasan et al. (2021) 44 BBC news Yes (HF, GitHub) ROUGE
@@
As shown in Table 1, GlotEval integrates publicly available multilingual benchmark datasets, covering machine translation, text classification, reading comprehension, summarization, and sequence labeling, spanning a wide range of languages from high-resource to low-resource. In total, GlotEval comprises 7 tasks and 22 benchmarks, evaluates in over 1500 languages, and utilizes 8 distinct metrics. Refer to Appendix A for more details of supported benchmark datasets.
@@ 3.2 Workflow
Once the user selects the desired benchmarks, languages, and prompt strategy, GlotEval’s data loader automatically locates each dataset, and filters and loads the relevant language subset(s) based on the user’s specification. It then initializes the appropriate model backend depending on the task type. Specifically, for non-generative tasks, we employ the HuggingFace Transformers backend (Wolf et al., 2020) to ensure more efficient use of computational resources. For generation tasks, such as machine translation, summarization, and open-ended text generation, we prioritize the vLLM backend (Kwon et al., 2023) to ensure high throughput, while retaining the HF Transformers generation interface for compatibility purposes.
After model inference is completed, GlotEval automatically computes evaluation metrics according to the task-specific settings listed in Table 1. Optionally, as mentioned before, by appending --store_details, users can export each sample’s prompt, model output, reference, and corresponding scores to a CSV file. This facilitates fine-grained error analysis and supports visual exploration of model behavior across languages and tasks.
3.3 A Deeper Look in Benchmark Data Loader
@@
Specifically, we process each benchmark-provided language code—which may appear in the form of ISO 639-1, 639-2/B (bibliographic), 639-2/T (terminological), ISO 639-3 codes, or even language names—by utilizing the iso639-lang  [^4] https://pypi.org/project/iso639-lang/ Python package. This allows us to retrieve all available mappings from the ISO 639-3 standard, including ISO 639-3 identifiers, ISO 639-2/B, 639-2/T, and ISO 639-1 codes. Using both exact and fuzzy matching strategies, we attempt to automatically identify the corresponding ISO 639-3 code for each language. A report is generated that documents, for each benchmark language, whether the match was exact or fuzzy, and whether it corresponds to an individual language or a macrolanguage in the ISO 639-3 standard.
@@
We further identify the script used by each dataset, using GlotScript  [^5] https://pypi.org/project/GlotScript/ library to detect the dominant script. Here we assume each dataset is primarily in one script. We randomly select up to 100 lines and attempt script recognition into ISO 15924 script code. This ensures each dataset obtains a <language>_<script> label, such as eng_Latn. The final ISO 639-3 code, along with the script code, is stored as the value in a language-to-code dictionary within the benchmark’s configuration file. Hence, each language + script combination is standardized in GlotEval for consistent usage across benchmarks.
Multilingual Prompt Builder
We constructed a dedicated command-line prompt builder to automatically prepare or adapt prompt templates for multilingual tasks. Figure 3(b) visualizes this process. In particular, the builder leverages Microsoft Translator to convert an instruction and/or few-shot prompt template from a given source language into 130+ target languages, while ensuring that placeholders (e.g., {src_text}) remain intact during translation.
These newly created multilingual prompts, are stored in the updated prompt library. As a result, each dataset’s prompts are aligned with the same <language>_<script> language taxonomy, enabling consistent, language-specific evaluation.
@@ 4.1 Efficiency Analysis
We benchmark GlotEval’s inference speed on six tasks: Flores200, Aya, and XLSum for generative tasks, and SIB-200, Global-MMLU, and WikiANN for non-generative tasks. All evaluations are conducted on 19 languages spanning diverse writing systems (e.g., Latin, Arabic, Cyrillic, Devanagari, Chinese, etc.). For each language, we sample 10 examples per task for evaluation. We choose Qwen2-1.5B model Yang et al. (2024) for evaluation. For generative tasks, we measure generation throughput (prefilling and decoding) with vLLM backend. For non-generative tasks, we measure classification throughput (prefilling only) using HF Transformers.
@@
For detailed throughput performance, Appendix B shows statistics on both GPU environments. They demonstrate that in general, Nvidia A100 consistently achieves higher throughput than AMD MI250X across both generative and non-generative tasks. Besides, this gap may also reflect the different backends between vLLM and HF Transformers. We further observe that scripts such as Devanagari or Amharic (amh_Ethi) often have lower throughput, potentially due to more complex tokenization. Lastly, summarization tasks like XLSum typically involve longer inputs and outputs than sentence-level translation tasks (e.g., Flores200), which increases the prefilling overhead and thus reduces the overall tokens/s.
@@ 4.2 Case Study on Multilingual Translation
A demonstration of prompt templates is shown in Table 2. For evaluation, we utilized NTREX128, a multi-aligned benchmark containing parallel texts across 128 languages, which is supported in GlotEval. In the multilingual prompting condition, we used built-in prompt builder in GlotEval, with the support of Microsoft Translator service, to automatically translate prompts into 134 languages supported by their platform. In our case study, 106 of these languages overlap with NTREX128 languages, allowing us to test performance across this diverse language set.
@@ 5 Conclusion and Future Work
In this work, we introduced GlotEval, a lightweight yet comprehensive framework for massively multilingual evaluation of large language models. By supporting consistent multilingual benchmarking, incorporating language-specific prompt templates, and supporting flexible non-English-centric translation setups, GlotEval enables reliable and fine-grained assessments of LLMs in diverse linguistic contexts—including low-resource settings often neglected by traditional benchmarks. Our case study on multilingual machine translation with two LLMs illustrates the utility of GlotEval in revealing the strengths and weaknesses of multilingual LLMs and in identifying directions for future optimization. Overall, GlotEval aims to encourage more inclusive, transparent, and holistic evaluations of language models across a wide array of languages and tasks, thereby advancing robust multilingual NLP research.
As for future work, we plan to explore and integrate more diverse and comprehensive multilingual benchmarks to better evaluate LLM performance. Plus, we will explore the integration of benchmarks that the synergistic combination of automatic and human evaluation; for example, this could achieved through our pilot development of a lightweight web interface that supports crowd-sourced and expert-driven evaluation to supplement the automatic evaluation [^6] Source code and documentation are available at https://github.com/MaLA-LM/GlotEval-HumanEval and https://gloteval-humaneval.readthedocs.io.
@@ Appendix B Throughput Statistics
Language Flores200(Eng-X) Aya XLSum SIB-200 Global-MMLU WikiANN
@@
[…]
Language Flores200(Eng-X) Aya XLSum SIB-200 Global-MMLU WikiANN
@@
[…]

### v2

@@ Abstract
Large language models (LLMs) are advancing at an unprecedented pace globally, with regions increasingly adopting these models for applications in their primary languages. Evaluating these models in diverse linguistic environments, especially in low-resource languages, has become a major challenge for academia and industry. Existing evaluation frameworks suffer from inconsistency across different benchmarks, being disproportionately focused on English and a handful of high-resource languages, thereby overlooking the realistic performance of LLMs in multilingual and lower-resource scenarios. To address this critical challenge of fragmented and inconsistent multilingual evaluation, we introduce GlotEval, a unified and lightweight framework that systematically integrates 27 benchmarks under a standardized ISO 639-3 language identifier system, allowing for seamless incorporation of new benchmarks. Supporting nine key tasks (machine translation, text classification, summarization, open-ended generation, reading comprehension, sequence labeling, intrinsic evaluation, instruction following and reasoning), spanning over dozens to hundreds of languages, GlotEval uniquely enables language-specific, cross-benchmark analysis and non-English-centric evaluations at a scale previously less practical for many researchers. This enables a precise diagnosis of model strengths and weaknesses in diverse linguistic contexts. A multilingual translation case study demonstrates GlotEval’s applicability for multilingual and language-specific evaluations.
GlotEval: github.com/MaLA-LM/GlotEval
@@ 1 Introduction
In recent years, driven by rapid progress in natural language processing and deep learning, large language models (LLMs) such as GPT-4 (OpenAI, 2023) and DeepSeek-R1 (DeepSeek-AI et al., 2025) have shown remarkable reasoning and generation capabilities across multiple languages and tasks. Although these models approach or surpass expert-level performance in certain high-resource languages (e.g., English), they often exhibit substantial performance fluctuations in other linguistic environments (Zhang et al., 2024). This discrepancy partially arises from the imbalance and scarcity of training data of low-resource languages, and partially from the limited multilingual coverage of current evaluation frameworks: many were originally designed for English or a few widely spoken languages, making it difficult to extend them efficiently to more diverse linguistic tasks or to adapt custom prompts and configurations for each language. Meanwhile, as LLMs proliferate worldwide and different regions rely on their respective local languages, large-scale (massively) multilingual evaluation involving numerous low-resource languages has emerged as a critical research direction.
@@
    Consistent Cross-benchmark Multilingual Evaluation. We integrate 27 existing multilingual benchmarks into a unified pipeline, by standardizing all ISO 639-3 language codes in the different benchmarks, [^1] https://iso639-3.sil.org/about which is an accepted standard with a good coverage of the world’s languages. By aligning benchmark language identifiers with ISO 639-3 codes, we enable evaluations for specific languages or language groups (e.g., Bantu, Dravidian, or Uralic languages), allowing the framework to automatically search among integrated benchmarks to find matching test sets. This mapping also makes it easier to incorporate new large-scale benchmarks that target mid- or low-resource languages, ensuring flexibility for future expansions.
@@
    Language-Specific Prompt Templates. Users can configure prompts for each language individually, thereby enabling more precise assessments of a model’s instruction-following ability across diverse linguistic settings. All templates are maintained in a centralized prompt library that supports multilingual benchmarks, allowing easy customization as needed. In this way, each task within a benchmark can be run potentially using prompts in the task’s original language, rather than defaulting to English prompts. To simplify cross-lingual adaptation, we also implemented Microsoft Translator integration that automatically propagates user-defined prompt templates from one single language to 130+ supported languages. [^2] https://learn.microsoft.com/en-us/azure/ai-services/translator/language-support
@@
    Non-English-Centered Machine Translation Evaluation. GlotEval is designed to break away from the traditional English-centric paradigm. Thanks to translation benchmarks featuring fully or partially multi-aligned datasets, GlotEval enables non-English-centered translation evaluations by allowing any supported language to serve as the pivot: users simply update the pivot language in the configuration to assess “any-to-pivot” / “pivot-to-any” translation directions. This flexibility ensures that GlotEval breaks from the traditional “English $\leftrightarrow$ other language” paradigm and adapts seamlessly to diverse, potentially low-resource, language pairs.
By bringing all these capabilities together in a cohesive framework, GlotEval aims to facilitate large-scale, in-depth evaluations of multilingual LLMs across both widely spoken and underrepresented languages, ultimately driving forward more inclusive LLM evaluation. Thus, GlotEval’s primary contribution is not the collection of new tasks, but the synergistic integration and standardization of existing benchmarks, which can be a robust tool for researchers and developers conducting massively multilingual LLM evaluation.
@@ 2 Related Work
Several evaluation toolkits and benchmarks have been developed to systematically assess LLMs. EleutherAI’s LM Evaluation Harness (Gao et al., 2023) is a widely adopted framework covering over 60 tasks, including multilingual datasets such as XNLI (15 languages) and Belebele (122 languages). UltraEval (He et al., 2024) improves modularity and supports FLORES-200 for multilingual translation. OpenAI Evals provides a highly flexible, community-driven framework, [^3] https://github.com/openai/evals and OpenCompass (Contributors, 2023) offers a comprehensive platform with broad support for datasets and models. MEGA (Ahuja et al., 2023) evaluates generative LLMs across diverse languages, with a focus on standard NLP benchmarks. LightEval (Fourrier et al., 2023) developed a flexible LLM evaluation framework that supports different backends.
@@ 3.1 Benchmarks, Languages and Metrics
Tatoeba Challenge v2023 Tiedemann (2020) 500+ Mixed short sents. Yes (GitHub) BLEU, ChrF
OpenSubtitles v2024 Lison and Tiedemann (2016) 93 Subtitles Yes (GitHub) BLEU, ChrF
@@
Summarization XLSum Hasan et al. (2021) 44 News Yes (HF, GitHub) ROUGE
MassiveSumm Long (Varab and Schluter, 2021) 55 News Yes (HF) ROUGE
MassiveSumm Short (Varab and Schluter, 2021) 88 News Yes (HF) ROUGE
Instruction Following BenchMAX Rule-based (Huang et al., 2025) 17 Verifiable instructions Yes (HF) Instruction-level Acc. etc.
Reasoning BenchMAX Math (Huang et al., 2025) 17 Grade School Math Yes (HF) Accuracy
BenchMAX Science (Huang et al., 2025) 17 Graduate-level Scientific QA Yes (HF) Accuracy
@@
As shown in Table 1, GlotEval integrates publicly available multilingual benchmark datasets, covering machine translation, text classification, summarization, open-ended generation, reading comprehension, sequence labeling, intrinsic evaluation, instruction following and reasoning, spanning a wide range of languages from high-resource to low-resource. In total, GlotEval comprises 9 tasks and 27 benchmarks, evaluates in over 1500 languages, and utilizes diverse metrics. Refer to Appendix A for more details of supported benchmark datasets.
@@ 3.2 Workflow
Upon selecting the desired benchmarks, languages, and prompt strategy, the user triggers GlotEval’s data loader to automatically locate each dataset and load the relevant language subsets. It then initializes the appropriate model backend depending on the task type. Specifically, for non-generative tasks, we employ the HuggingFace Transformers backend (Wolf et al., 2020) to ensure more efficient use of computational resources. For generation tasks, such as machine translation, summarization, and open-ended text generation, we prioritize the vLLM backend (Kwon et al., 2023) to ensure high throughput, while retaining the HF Transformers generation interface for compatibility purposes.
After model inference is completed, GlotEval automatically computes evaluation metrics according to the task-specific settings listed in Table 1. Optionally, as mentioned before, by appending --store_details, users can export each sample’s prompt, model output, reference, and corresponding scores to a JSONL file, which allows researchers to work outside the framework and conduct custom error analysis and result visualization. This ensures that our framework is not just an evaluation executor, but also a starting point for more fine-grained analysis.
3.3 A Deeper Look at Benchmark Data Loader
@@
Specifically, we process each benchmark-provided language code—which may appear in the form of ISO 639-1, 639-2/B (bibliographic), 639-2/T (terminological), ISO 639-3 codes, or even language names—by utilizing the iso639-lang Python package. [^4] https://pypi.org/project/iso639-lang/ This allows us to retrieve all available mappings from the ISO 639-3 standard, including ISO 639-3 identifiers, ISO 639-2/B, 639-2/T, and ISO 639-1 codes. Using both exact and fuzzy matching strategies, we attempt to automatically identify the corresponding ISO 639-3 code for each language. A report is generated that documents, for each benchmark language, whether the match was exact or fuzzy, and whether it corresponds to an individual language or a macrolanguage in the ISO 639-3 standard.
We further identify the script used by each dataset, using GlotScript (Kargaran et al., 2024) to detect the dominant script. [^5] https://pypi.org/project/GlotScript/ Here we assume each dataset is primarily in one script. We randomly select up to 100 lines and attempt script recognition into ISO 15924 script code. This ensures each dataset obtains a <language>_<script> label, such as eng_Latn. The final ISO 639-3 code, along with the script code, is stored as the value in a language-to-code dictionary within the benchmark’s configuration file. Hence, each language + script combination is standardized in GlotEval for consistent usage across benchmarks.
Multilingual Prompt Builder
We constructed a dedicated command-line prompt builder to automatically prepare or adapt prompt templates for multilingual tasks. Figure 3(b) visualizes this process. In particular, the builder leverages Microsoft Translator to convert an instruction and/or few-shot prompt template from a given source language into 130+ target languages, while ensuring that placeholders (e.g., {src_text}) remain intact during translation. These newly created multilingual prompts, are stored in the updated prompt library. As a result, each dataset’s prompts are aligned with the same <language>_<script> language taxonomy, enabling consistent, language-specific evaluation.
Note that the automatic translation of prompts is intended as a convenience feature to support rapid, large-scale multilingual evaluation. While translation quality may vary—particularly for low-resource languages—this approach offers a practical starting point for exploratory analysis with language-specific prompts at scale. The framework remains fully customizable: users are able to provide their own human-written or verified prompts in the prompt library for languages of interest.
@@ 4.1 Efficiency Analysis
We benchmark GlotEval’s inference speed on six tasks: FLORES-200, Aya, and XLSum for generative tasks, and SIB-200, Global-MMLU, and WikiANN for non-generative tasks. All evaluations are conducted on 19 languages spanning diverse writing systems (e.g., Latin, Arabic, Cyrillic, Devanagari, Chinese, etc.). For each language, we sample 10 examples per task for evaluation. We choose Qwen2-1.5B model Yang et al. (2024) for evaluation. For generative tasks, we measure generation throughput (prefilling and decoding) with vLLM backend. For non-generative tasks, we measure classification throughput (prefilling only) using HF Transformers.
@@
For detailed throughput performance, Appendix B shows statistics on both GPU environments. They demonstrate that in general, NVIDIA A100 consistently achieves higher throughput than AMD MI250X across both generative and non-generative tasks. Besides, this gap may also reflect the different backends between vLLM and HF Transformers. We further observe that scripts such as Devanagari or Amharic (amh_Ethi) often have lower throughput, potentially due to more complex tokenization. Lastly, summarization tasks like XLSum typically involve longer inputs and outputs than sentence-level translation tasks (e.g., FLORES-200), which increases the prefilling overhead and thus reduces the overall tokens/s.
@@ 4.2 Case Study on Multilingual Translation
A demonstration of prompt templates is shown in Table 2. For evaluation, we utilized NTREX-128, a multi-aligned benchmark containing parallel texts across 128 languages, which is supported in GlotEval. In the multilingual prompting condition, we used built-in prompt builder in GlotEval, with the support of Microsoft Translator service, to automatically translate prompts into 134 languages supported by their platform. In our case study, 106 of these languages overlap with NTREX-128 languages, allowing us to test performance across this diverse language set.
@@ 5 Conclusion and Future Work
In this work, we introduced GlotEval, a lightweight yet comprehensive framework for massively multilingual evaluation of LLMs. By supporting consistent multilingual benchmarking, incorporating language-specific prompt templates, and supporting flexible non-English-centric translation setups, GlotEval enables consistent assessments of LLMs in diverse linguistic contexts—including low-resource settings often neglected by traditional benchmarks. Our case study on multilingual machine translation with two LLMs illustrates the utility of GlotEval in revealing the strengths and weaknesses of multilingual LLMs and in identifying directions for future optimization. Overall, GlotEval aims to encourage more inclusive, transparent, and holistic evaluations of language models across a wide array of languages and tasks, thereby advancing robust multilingual NLP research.
As for future work, we plan to integrate more diverse and comprehensive multilingual benchmarks to better evaluate LLM performance. Plus, we will explore the integration of benchmarks that the synergistic combination of automatic and human evaluation; for example, this could achieved through our pilot development of a lightweight web interface that supports crowd-sourced and expert-driven evaluation to supplement the automatic evaluation. [^6] Source code and documentation are available at https://github.com/MaLA-LM/GlotEval-HumanEval and https://gloteval-humaneval.readthedocs.io
@@ Acknowledgments
Sawal Devkota, Bhavani Sai Praneeth Varma Mantina, Ananda Sreenidhi, Mengjie Wang, and Samea Yusofi contributed to this project as part of the “Data Analysis Software Project for Natural Language” course at TU Darmstadt, under the guidance of Shaoxiong Ji. This teaching activity was funded by LOEWE Center DYNAMIC as part of the Hessian program for the promotion of cutting-edge research LOEWE under the grant number of LOEWE1/16/519/03/09.001(0009)/98.
@@ Appendix B Throughput Statistics
Language FLORES-200(Eng-X) Aya XLSum SIB-200 Global-MMLU WikiANN
@@
Table 3: Throughput with NVIDIA A100 40GB GPU. Each cell contains: $\frac{\#\text{generated tokens}}{\text{wall time (seconds)}}=\text{average tokens/s}$.
Language FLORES-200(Eng-X) Aya XLSum SIB-200 Global-MMLU WikiANN
@@
Table 4: Throughput with AMD MI250X 64GB GPU. Each cell contains: $\frac{\#\text{generated tokens}}{\text{wall time (seconds)}}=\text{average tokens/s}$.